OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME? (if you are organized and good at explaining) 2(x - 1) ≤ 12

OpenStudy (nikato):

Solve for x?

OpenStudy (nikato):

First divide both sides by 2. What do you get?

OpenStudy (anonymous):

you have to take two equations... 1. 2(x - 1) = 12 2. 2(x - 1) < 12 so, do the same as Mr. nikato said, you will get two number pointing x =... and x<... then it will be (...,...)

OpenStudy (anonymous):

\[2(x-1)\leq 12\] divide both sides by \(2\) to cancel the 2 on the left \[x-1\leq 12\div 2\\ x-1\leq 6\]

OpenStudy (anonymous):

then finish by adding 1 to both sides to get \(x\) by itself \[x-1+1\leq 6+1\\ x\leq 7\]